# Entanglement scaling and charge fluctuations in a Fermi liquid of composite fermions

### v2

@@ I Introduction
Here we report a systematic study of entanglement in several CFLs described by the Rezayi-Read wave function [26], known to have a large overlap with the Coulomb ground state [29, 30]. We assume full spin polarization and compare the fermionic CFLs at fillings $\nu=1/2$ and $\nu=1/4$, which are distinguished by the number of vortices attached to each electron (two and four, respectively), and contrast those against an odd number of vortices in bosonic CFLs at $\nu=1$ and $\nu=1/3$ fillings. We employ Monte Carlo techniques to evaluate the second Rényi entropy using the SWAP algorithm [31]. In all CFL states considered, we find a pronounced enhancement of entanglement scaling, consistent with Ref. [27], up to the largest system sizes accessible ($N\lesssim 60$). At the same time, we reconcile these findings with entropy decompositions considered in Ref. [28] by showing that the latter is more sensitive to finite-size effects and LLL projection compared to the total entropy. We demonstrate the robustness of our conclusions in sphere and torus geometries. Finally, we use the same setup to evaluate charge fluctuations in a subsystem, for which analytical predictions are available [32]. We confirm that charge fluctuations in the CFL state follow the expected area-law scaling, in sharp contrast with free fermions. Moreover, within the same range of system sizes, we accurately extract the subleading corner contribution to the charge fluctuations, in agreement with the expected “super-universal” value [32] that also holds for gapped [33, 34] and critical states [35, 36] in (2+1)-dimension.
@@ II The CFL wave function and its Rényi entropy
In the asymptotic limit, for free fermions in 2D, the area-law scaling of $S_{2}$ is violated by a term that depends logarithmically on the dimensionless quantity $\lambda\equiv k_{F}r_{A}$, with $k_{F}$ denoting the Fermi momentum. The violation is given by the Widom formula [19, 20, 21]:
@@
where $\kappa_{\mathrm{w}}$ is the Widom coefficient that depends solely on the geometry and effective central charge of a (1+1)-dimensional chiral relativistic fermion [46, 47]. For a circular 2D Fermi surface, we have $\kappa_{\mathrm{w}}=1/4$. One of the goals of this paper is to ascertain whether the entropy of CFL states follows the scaling $S_{2}=\kappa\lambda\ln\lambda+\ldots$, analogous to Eq. (4), and whether $\kappa$ is the same as $\kappa_{\mathrm{w}}$. In the remainder of this section, we directly evaluate the $S_{2}$ Rényi entropy for the wave functions in Eq. (1) on the torus and sphere, and compare them against Eq. (4).
@@ II.1 Torus
In Fig. 1(a)-(b) we show the evaluated $S_{2}$ for bosonic and fermionic CFL states at fillings $\nu=1/m$, with $m{=}1,2,3,4$. Based on our free-fermion benchmarks presented in Appendix D, to minimize the finite-size effect, we harvest data points such that the subsystem area ranges from 1% to 30% of the total area and use a fixed Fermi momentum of $k_{F}=\sqrt{2\nu}$. The extrapolated slope $\kappa$, in units of the Widom slope $\kappa_{\mathrm{W}}$, is plotted in Fig. 1(c) as a function of inverse system size.
@@ II.2 Sphere
where $N_{0,L,M}$ is a normalization factor, and the flux for the physical electrons is $2Q=2(N{-}1)$. Given that we only have one power of Jastrow, the derivatives can be evaluated as follows [9]:
@@
The above construction produces CFL states at general fillings. For the fermionic state at $\nu=1/4$, we choose to only project with two Jastrow factors, leaving the other two outside – this is similar to the $\bm{\alpha}=(4,0)$ state on the torus. For the bosonic states at $\nu=1$ and $\nu=1/3$, we need to divide or multiply by a single Jastrow factor respectively [55, 30], again mirroring the construction on the torus. It is known that the microscopic CFL wave function is not very sensitive to the number of Jastrows placed inside the LLL-projection operator [56, 57]. Note that $\nu=1$ is a special case as as the hardcore constraint is not enforced properly, which can result in instability after projection. To fix this, we implicitly enforce a non-zero hardcore radius in our Monte Carlo simulation.
@@
Fig. 2 demonstrates the similarity of CFL results on the sphere with those presented previously for the torus geometry. This consistency highlights the robustness of our results. Furthermore, for the sequence of system sizes $N=n^{2}$ we consider on the sphere, the CFL has a uniform angular momentum $L=0$ state, which removes some of the ambiguity associated with the definition of the Fermi momentum on the torus.
@@ III Charge fluctuations
For example, in non-interacting 2D Fermi liquids, $S_{2}$ and $\Delta^{2}N_{A}$ differ only by a multiplicative constant [58, 59]. Similarly, in a NFL, charge fluctuations are expected to closely track entanglement entropy [60]. Recently, the scaling of charge cumulants or “disorder operators” was indeed shown to exhibit a logarithmic violation of the area law in NFLs at a quantum critical point [61]. Since $\Delta^{2}N_{A}$ can be computed with a similar Monte Carlo method and the same system sizes as $S_{2}$, we can leverage the better analytical understanding of charge fluctuations to support our previous entanglement results.
Remarkably, Fig. 3 shows that $\Delta^{2}N_{A}$ in CFL states, with and without LLL projection, obeys an area law – in stark contrast with ordinary Fermi liquids. This can be understood from the fact that the electrons in the CFL exhibit an effectively finite correlation length, leading to an exponentially suppressed tunneling density of states [62, 63, 64, 65] and a Coulomb gap observed in experiment [66, 67, 68, 69, 70]. This implies that charge fluctuations of the physical electrons forming the CFL are gapped and follow an area law [32, 71]. This striking property of the microscopic CFL wave function, to the best of our knowledge, has been demonstrated for the first time in this work.
@@ IV Modulus and sign structure
$S_{2}(\Psi)=S_{2,\mathrm{p}}+S_{2,\mathrm{mod}}++S_{2,\mathrm{sign}}.$ (14)
@@
The hope behind the decomposition of Eq. (14) is that the sign structure of the wave function carries the leading contribution to the entanglement as $N{\to}\infty$. While there are examples of sign wave functions with provable large entanglement [74, 75], there are also instances where the mod entropy is not dominated by that of the sign [76]. Thus, the scaling of different terms with $N$ in Eq. (14) is not rigorously understood in general. In fact, even for free fermions, the scalings of $S_{2}(|\Psi|)$ and $S_{2,\mathrm{sign}}$ are not known analytically. Our Monte Carlo analysis for free fermions presented below is consistent with $S_{2,\mathrm{sign}}$ carrying the dominant $\lambda\ln\lambda$ dependence, however both $S_{2}(|\Psi|)$ and $S_{2,\mathrm{sign}}$ suffer from much more pronounced finite-size effects compared to the total entropy $S_{2}$. Consequently, we will argue that the decomposition in Eq. (14) obscures the interpretation of the data as it replaces a tractable free-fermion problem with an effectively “many-body” object that carries non-trivial finite-size dependence due to the mixing of subleading [$\mathcal{O}(\lambda)$ and $\mathcal{O}(\ln\lambda)$] terms. Below we discuss in detail the entropy decompositions for free fermions and then apply the same considerations to the CFL case.
@@ IV.1 Mod and sign entropy for free fermions
The scaling of different contributions to the entropy in Eq. (14) is not rigorously known, even for free fermions where the full entropy has an established analytic expression. Fig. 4(a) contrasts the full $S_{2}$ entropy for free fermions against its decompositions in Eq. (14) obtained via Monte Carlo, Fig. 4(b)-(d). While the full entropy can be computed for large free-fermion systems containing thousands of particles, the evaluation of mod-sign entropy contributions, to the best of our knowledge, can only be done with Monte Carlo. This limits the accessible system sizes to $N\lesssim 100$, which is similar to the range for the CFL states.
Within the range $N\lesssim 100$, we observe the full entropy $S_{2}$ steadily converging to the Widom slope in Fig. 4(a). Figs. 4(b)-(c) illustrate the importance of considering the $\mathrm{p}$ and $\mathrm{mod}$ contributions to the entropy separately. They each appear to converge nicely to the respective scalings, $S_{2,p}\sim\mathcal{O}(\ln\lambda)$ and $S_{2,\mathrm{mod}}\sim\mathcal{O}(\lambda)$. However, if we add them together to make up $S_{2}(|\Psi|)$, Fig. 4(c), the latter converges non-monotonically when plotted as $S_{2}(|\Psi|)/\lambda$ vs $\ln\lambda$. If we estimate the slope of $S_{2}(|\Psi|)$ by adding the fits, we obtain the solid line in Fig. 4(c), which suggests that our system sizes are not large enough to see the expected area-law scaling $S_{2}(|\Psi|)\sim\lambda$ in the thermodynamic limit. This has an adverse impact on the scaling of the sign entropy in Fig. 4(d), which overshoots the Widom slope to compensate for the downturn in $S_{2}(|\Psi|)$. By plotting the Widom slope minus the extrapolation of $S_{2}(|\Psi|)$ using a solid line in Fig. 4(d), we see that it might only approach the Widom scaling in much larger system sizes ($\ln\lambda\gtrsim 4$). At the same time, the full entropy in Fig. 4(a) tracks the Widom scaling much more closely even in smaller system sizes. As the range of system sizes in this example is comparable to the CFL case, the usefulness of mod and sign decompositions over the full entropy is questionable.
@@ Appendix E The Widom formula on the sphere
with a discrete sum over these modes. The flux factor $|\hat{n}_{x}\cdot\hat{n}_{k}|/2\pi$ needs to be replaced by $1/2\pi$, accounting for the rotational symmetry of the cap around the axis, and the double counting factor of $1/2$ becomes unnecessary. Then, the Rényi entropy for a spherical cap of angle $\theta$ is:
@@
$=$ $\frac{c_{\mathrm{eff}}(1+\alpha)}{12\alpha}(2\sqrt{N}-1)\sin\theta\ln(R\sin\theta),$
To enable a comparison between this and CFL states, we need to control the particle density through the flux. This, in turn, affects the definition of the radius $R$. For the CFL, the radius of the sphere is $R=\sqrt{Q}$ and the maximum occupied effective angular momentum is $L_{\mathrm{max}}=\sqrt{N}$. Setting the radius of the cap to $R\sin\theta=\sqrt{Q}\sin\theta$, and the effective Fermi wave number $k_{F}{=}L_{\mathrm{max}}/R=\sqrt{N/Q}$, we express the entropy in terms of the dimensionless length $\lambda_{s}{=}k_{F}\sqrt{Q}\sin\theta$ :

### v3

@@ I Introduction
Here we report a systematic study of entanglement in several CFLs described by the Rezayi-Read wave function [26], known to have a large overlap with the LLL Coulomb ground state [29, 30]. We assume full spin polarization and compare the fermionic CFLs at fillings $\nu=1/2$ and $\nu=1/4$, which are distinguished by the number of vortices attached to each electron (two and four, respectively), and contrast those against an odd number of vortices in bosonic CFLs at $\nu=1$ and $\nu=1/3$ fillings. We employ Monte Carlo techniques to evaluate the second Rényi entropy using the SWAP algorithm [31]. In all CFL states considered, we find a pronounced enhancement of entanglement scaling, consistent with Ref. [27], up to the largest system sizes accessible ($N\lesssim 60$). At the same time, we reconcile these findings with entropy decompositions considered in Ref. [28] by showing that the latter is more sensitive to finite-size effects and LLL projection compared to the total entropy. We demonstrate the robustness of our conclusions in sphere and torus geometries. Finally, we use the same setup to evaluate charge fluctuations in a subsystem, for which analytical predictions are available [32]. We confirm that charge fluctuations in the CFL state follow the expected area-law scaling, in sharp contrast with free fermions. Moreover, within the same range of system sizes, we accurately extract the subleading corner contribution to the charge fluctuations, in agreement with the expected “super-universal” value [32] that also holds for gapped [33, 34] and critical states [35, 36] in (2+1)-dimension.
@@ II The CFL wave function and its Rényi entropy
In the asymptotic limit, for free fermions in 2D, the area-law scaling of $S_{2}$ is violated by a term that depends log-linearly on the dimensionless quantity $\lambda\equiv k_{F}r_{A}$, with $k_{F}$ denoting the Fermi momentum. The violation is given by the Widom formula [19, 20, 21]:
@@
where $\kappa_{\mathrm{w}}$ is the Widom coefficient that depends solely on the geometry and effective central charge of a (1+1)-dimensional chiral relativistic fermion [46, 47]. For a circular 2D Fermi surface, we have $\kappa_{\mathrm{w}}=1/4$. One of the goals of this paper is to ascertain whether the entropy of CFL states follows the scaling $S_{2}=\kappa\lambda\ln\lambda+\ldots$, analogous to Eq. (4) and whether $\kappa$ is the same as $\kappa_{\mathrm{w}}$. In the remainder of this section, we directly evaluate the $S_{2}$ Rényi entropy for the wave functions in Eq. (1) on the torus and sphere, and compare them against Eq. (4).
@@ II.1 Torus
In Fig. 1(a)-(b) we show the evaluated $S_{2}$ for bosonic and fermionic CFL states at fillings $\nu=1/m$, with $m{=}1,2,3,4$. Based on our free-fermion benchmarks presented in Appendix D, to minimize the finite-size effect, we harvest data points such that the subsystem area ranges from 1% to 30% of the total area and use a fixed Fermi momentum of $k_{F}=\sqrt{2\nu}/\ell$. The extrapolated slope $\kappa$, in units of the Widom slope $\kappa_{\mathrm{W}}$, is plotted in Fig. 1(c) as a function of inverse system size.
@@ II.2 Sphere
where $N_{0,L,M}$ is a normalization factor, and the flux for the physical electrons is $2Q=2(N{-}1)$. Given that we only have one power of Jastrow factors, the derivatives can be evaluated as follows [9]:
@@
The above construction produces CFL states at general fillings. For the fermionic state at $\nu=1/4$, we choose to only project with two Jastrow factors, leaving the other two outside – this is similar to the $\bm{\alpha}=(4,0)$ state on the torus. For the bosonic states at $\nu=1$ and $\nu=1/3$, we need to divide or multiply by a single Jastrow factor respectively [55, 30], again mirroring the construction on the torus. It is known that the microscopic CFL wave function is not very sensitive to the number of Jastrows placed inside the LLL-projection operator [56, 57]. Note that $\nu=1$ is a special case as the hardcore constraint is not enforced properly, which can result in instability after projection. To fix this, we implicitly enforce a non-zero hardcore radius in our Monte Carlo simulation.
@@
Fig. 2 demonstrates the similarity of CFL results on the sphere with those presented previously for the torus geometry. This consistency highlights the robustness of our results. Furthermore, for the sequence of system sizes $N=n^{2}$ we consider on the sphere, the CFL is a uniform angular momentum $L=0$ state, which removes some of the ambiguity associated with the definition of the Fermi momentum on the torus.
@@ III Charge fluctuations
For example, in non-interacting 2D Fermi liquids, $S_{2}$ and $\Delta^{2}N_{A}$ differ only by a multiplicative constant [58, 59]. Similarly, in a NFL, charge fluctuations are expected to closely track entanglement entropy [22]. Recently, the scaling of charge cumulants or “disorder operators” was indeed shown to exhibit a logarithmic violation of the area law in NFLs at a quantum critical point [60]. Since $\Delta^{2}N_{A}$ can be computed with a similar Monte Carlo method for the same system sizes as $S_{2}$, we can leverage the better analytical understanding of charge fluctuations to support our previous entanglement results.
Remarkably, Fig. 3 shows that $\Delta^{2}N_{A}$ in CFL states, with and without LLL projection, obeys an area law. This implies that charge fluctuations of the physical electrons forming the CFL are gapped [32, 61], in stark contrast with ordinary Fermi liquids. While it is well-known that the $\nu{=}1/2$ CFL state does not exhibit a quantized Hall resistance [62], transport experiments probe a gap to creating or removing fractionalized excitations. By contrast, $\Delta^{2}N_{A}$ probes the fluctuation of electrons in a subsystem and is therefore more similar to scanning tunneling spectroscopy experiments [63, 64, 65, 66, 67], which indeed observe a Coulomb gap at $\nu{=}1/2$. This is because the electrons in the CFL exhibit an effectively finite correlation length leading to an exponentially suppressed tunneling density of states [68, 69, 70, 71]. The area-law for $\Delta^{2}N_{A}$ obeyed by the microscopic CFL wave function, to the best of our knowledge, has been demonstrated for the first time in this work.
@@ IV Modulus and sign structure
$S_{2}(\Psi)=S_{2,\mathrm{p}}+S_{2,\mathrm{mod}}+S_{2,\mathrm{sign}}.$ (14)
@@
The hope behind the decomposition of Eq. (14) is that the sign structure of the wave function carries the leading contribution to the entanglement as $N{\to}\infty$. While there are examples of wave functions with provable large sign entanglement [74, 75], there are also instances where the mod entropy is not dominated by that of the sign [76]. Thus, the scaling of different terms with $N$ in Eq. (14) is not rigorously understood in general. In fact, even for free fermions, the scalings of $S_{2}(|\Psi|)$ and $S_{2,\mathrm{sign}}$ are not known analytically. Intuitively, the reason for this is that the decomposition in Eq. (14) replaces a tractable free-fermion problem with an effectively “many-body” object that no longer has a simple description in terms of correlation matrices [43]. Moreover, $S_{2}(|\Psi|)$ and $S_{2,\mathrm{sign}}$ carry non-trivial finite-size dependence due to the mixing of subleading [$\mathcal{O}(\lambda)$ and $\mathcal{O}(\ln\lambda)$] terms. Our Monte Carlo analysis for free fermions presented below is consistent with $S_{2,\mathrm{sign}}$ carrying the dominant $\lambda\ln\lambda$ dependence, however both $S_{2}(|\Psi|)$ and $S_{2,\mathrm{sign}}$ suffer from much more pronounced finite-size effects compared to the total entropy $S_{2}$. Below we discuss in detail the entropy decompositions for free fermions and then apply the same considerations to the CFL case.
@@ IV.1 Mod and sign entropy for free fermions
As mentioned previously, the scaling of different contributions to the entropy in Eq. (14) is not rigorously known, even for free fermions where the full entropy has an established analytic expression. Fig. 4(a) contrasts the full $S_{2}$ entropy for free fermions against its decompositions in Eq. (14) obtained via Monte Carlo, Fig. 4(b)-(d). While the full entropy can be computed for large free-fermion systems containing thousands of particles, the evaluation of mod-sign entropy contributions, to the best of our knowledge, can only be done with Monte Carlo. This limits the accessible system sizes to $N\lesssim 100$, which is similar to the range for the CFL states.
Within the range $N\lesssim 100$, we observe the full entropy $S_{2}$ steadily converging to the Widom slope in Fig. 4(a). Figs. 4(b)-(c) illustrate the importance of considering the $\mathrm{p}$ and $\mathrm{mod}$ contributions to the entropy separately. They each appear to converge nicely to the respective scalings, $S_{2,p}\sim\mathcal{O}(\ln\lambda)$ and $S_{2,\mathrm{mod}}\sim\mathcal{O}(\lambda)$. However, if we add them together to make up $S_{2}(|\Psi|)$, Fig. 4(c), the latter converges non-monotonically when plotted as $S_{2}(|\Psi|)/\lambda$ vs $\ln\lambda$. If we estimate the slope of $S_{2}(|\Psi|)$ by adding the fits, we obtain the solid line in Fig. 4(c), which suggests that our system sizes are not large enough to see the expected area-law scaling $S_{2}(|\Psi|)\sim\lambda$ in the thermodynamic limit. This adversely impacts the scaling of the sign entropy in Fig. 4(d), which overshoots the Widom slope to compensate for the downturn in $S_{2}(|\Psi|)$. By plotting the Widom slope minus the extrapolation of $S_{2}(|\Psi|)$ using a solid line in Fig. 4(d), we see that it might only approach the Widom scaling in much larger system sizes ($\ln\lambda\gtrsim 4$). At the same time, the full entropy in Fig. 4(a) tracks the Widom scaling much more closely even in smaller system sizes. As the range of system sizes in this example is comparable to the CFL case, the usefulness of mod and sign decompositions over the full entropy is questionable.
@@ Appendix E The Widom formula on the sphere
with a discrete sum over these modes. Here, $c_{\mathrm{eff}}$ denotes the effective central charge of a (1+1)- dimensional chiral relativistic fermion, the vectors $\hat{\mathbf{n}}_{x}$ and $\hat{\mathbf{n}}_{k}$ are unit normals for real space boundary and the Fermi surface, respectively, and $\alpha$ denotes the Rényi index (we will mainly focus on $\alpha=2$).
The flux factor $|\hat{n}_{x}\cdot\hat{n}_{k}|/2\pi$ needs to be replaced by $1/2\pi$, accounting for the rotational symmetry of the cap around the axis, and the double counting factor of $1/2$ becomes unnecessary. Then, the Rényi entropy for a spherical cap of angle $\theta$ is:
@@
$=$ $\frac{c_{\mathrm{eff}}(1+\alpha)}{12\alpha}(2\sqrt{N}-1)\sin\theta\ln(R\sin\theta).$
To enable a comparison between this and CFL states, we need to control the particle density through the flux. This, in turn, affects the definition of the radius $R$. For the CFL, the radius of the sphere is $R=\sqrt{Q}$ and the maximum occupied effective angular momentum is $L_{\mathrm{max}}=\sqrt{N}$. Setting $\alpha=2$, the radius of the cap to $R\sin\theta=\sqrt{Q}\sin\theta$, and the effective Fermi wave number $k_{F}{=}L_{\mathrm{max}}/R=\sqrt{N/Q}$, we express the entropy in terms of the dimensionless length $\lambda_{s}{=}k_{F}\sqrt{Q}\sin\theta$ :